# GANsemble for Small and Imbalanced Data Sets:A Baseline for Synthetic Microplastics Data

### v1

@@ Abstract.
*daniel.platnick@torontomu.ca
@@ 1. Introduction
Oversampling the minority class is a natural remedy for class imbalance [Oversample]. Synthetic Minority Oversampling Technique (SMOTE) is a well known method which produces synthetic samples to fix class imbalances using K-nearest neighbors in the feature space [SMOTE]. [heart_spectral_aug] [heart_spectral_aug] performed analysis on different augmentation strategies for convolutional neural network (CNN) classification of medical image spectra, and found that simple strategies such as masking and horizontal flipping can greatly improve model performance. Today, oversampling techniques typically include intelligent data augmentations [augmentation_survey]. [CutMix] [CutMix] introduced a state-of-the-art image augmentation algorithm called CutMix, which maintains the benefits of regional dropout regularization without loss of information. [no_best_resampling_approach] [no_best_resampling_approach] showed that cases may vary depending on the nature of the data. Therefore, given imperfect information, the system designer must decide the best augmentation strategy to use. This is the challenge of augmentation selection.
@@
Inspired by recent developments in generative AI [GAN, cGAN], we present GANsemble to overcome this challenge and further microplastics research. The proposed method connects ideas from data augmentation and synthetic data generation [cGAN_synthdata] to provide automatic augmentation strategy selection and enhance model learning on small and imbalanced data sets. The GANsemble data chooser module enables the AI system to manipulate the inputted data through augmentation, generating new augmentation strategies to find the strategy which maximizes model generalization. Specifically, GANsemble performs an n-step factorial search search on inputted augmentation strategies, and employs a deep neural network (DNN) to intelligently select the best strategy $Aug^{*}$. This strategy is then used to oversample the original data set and train a cGAN to generate new synthetic samples. GANsemble can significantly improve the augmentation selection process, by decreasing the need for human intervention. Furthermore, we provide MPcGAN, the first cGAN for generating synthetic microplastics (SYMP) data, and establish baseline FID and IS scores for SYMP data.
@@
Generating synthetic sample data to enhance the training set is an idea increasing in popularity [cGAN, diffusion_sample, diffusion_sample, CVAE]. This approach has not been applied to microplastics research. To make progress in SYMP data generation, this work applies cGANs to create SYMP data, and baseline SYMP data quality measures are established. MPcGAN is trained to produce class-conditioned SYMP samples that enhance training set size, balance, and variation. Note, this paper builds on the important work of [source_dataset] [source_dataset], which identified the most effective amount of oversampling with augmentation for a small imbalanced microplastics data set. However, Tian et al. introduce data leakage into their analysis by creating augmented data from the training set, and evaluating on the same training set. This is a problem, because in practice, the model will be evaluated on data it has never seen before, not data created from previous training data. Performing a similar study, this research builds on their analysis by using a separate evaluation set to prevent data leakage between train and test sets. Multiple runs are used to increase experimental validity.
@@ 2. Data Set Description
The data set we use was introduced by [source_dataset] [source_dataset], and is the only easily accessible microplastics data set we could find. The raw data and supplementary information on data preparation can be found from these sources [source_dataset, data_source2, data_source3]. In this section, we review its characteristics and how it was created.
Figure 1. Example of the polar conversion that allows DNNs to treat polymer spectra as images [source_dataset].
[…]
Tian et al. [source_dataset] also introduced a visual polarized coordinate transformation which allows microplastic particle spectra to be treated as as images. This enables the use of computer vision models, which have been found to be very useful in practice. The spectra in the microplastics data set are mapped through the polar conversion process as follows: The wavenumbers 975-1800 are mapped from 0°$\degree$ to 360°$\degree$ and the absorption rate becomes the polar coordinate magnitude. The images are then colourized such that colour intensity grows with the absorption value. Figure 1 shows an example of this transformation visually.
@@ 3.1. The Data Chooser Module
The second input is a list of base augmentation strategies to be used. While any set of base data augmentation strategies can be used, for the microplastic data, we use four Gaussian approaches: (1) horizontal flipping and horizontal shifting, (2) blur and rotation, (3) zoom and rotation, and (4) circular and rectangular masking. Base augmentation strategy 1 performs random horizontal flipping and shifting. Random horizontal flips in spectra data have been shown to greatly improve ML classification performance [heart_spectral_aug], while shifting ensures there are not too many duplicate samples. Augmentation strategy 2 uses varying levels of blur and random rotations to distort the spectra images. The blur effect is achieved through a Gaussian filter with $\alpha\in[10,23]$ and $\sigma\in[2.8,3.82]$, where $\alpha$ is blur intensity and $\sigma$ is smoothness. Different $\alpha$ and $\sigma$ values were tried and these ranges gave the best results. Augmentation strategy 3 applies random rotations and zoom between $1.0$x and $1.34$x. Augmentation strategy 4 is a Monte Carlo style masking strategy, where circles and rectangles cover different areas of the polymer spectra. Masking has also been effective for augmenting data in heart spectra classification [heart_spectral_aug], as it forces the learning algorithm to focus on different regions of the image when masked regions are not present.
@@
In principle, it would be possible to train all $\mathcal{O}(2^{A})$ possible cGANs and thereby directly evaluate the effectiveness of the different combinations. However, even for a small data set, doing so is prohibitively expensive given the time and lack of stability of GAN training. As such, we instead focus on the simpler task of classification using a pre-trained model. That is, we consider each of the $\mathcal{O}(2^{n})$ possible combinations of augmentation methods, and train a classifier $r$ times each, where $r$ is another hyperparameter controlling how extensively the augmentation strategies are explored. The classifier with the best average performance over the $r$ runs is then chosen and used for training the cGAN. For the microplastics data, we use a fully pre-trained ResNet50 as the classifier for identifying the best augmentation strategy. We refer to the resulting augmentation strategy as $Aug^{*}$.
@@ 4. Experiments and Results
Recall that our data set contains only 210 images. This was split into an unbiased testing set containing 20 samples, with 2 images per class, and a training set containing the remaining 190 samples. The 20 samples were not used in creating the augmented data sets from Table 1 or Figure 2 to prevent data leakage. The 20 samples were later added to help train the cGAN. All code relating to this work can be found here: https://github.com/DanielPlatnick/GANsemble
@@ 4.1. Identifying An Effective Size For Augmented Data Sets
In our first experiment, we attempt to reproduce the experiments by Tian et al. [source_dataset], in which they evaluated the tradeoff between accuracy and training time related to augmented data set size. In the original work, a custom augmentation strategy was used. Our goal is to evaluate if similar conclusions can be made with simpler augmentation strategies and without the data leakage that occurred in the original work.
For simplicity, we tested with horizontal flipping and horizontal shifting. Using this augmentation strategy, we create 10 purely augmented data sets of sizes $100,200,\ldots,1000$, each with balanced classes. We then used this data to train three different DNN models: a CNN with $13,804,510$ trainable parameters, a ResNet50 with $23,528,522$ trainable parameters, and a version of that ResNet50 pre-trained on ImageNet [Resnet, ImageNet]. Each DNN was trained 5 times and the average results on our test set of 20 images are shown in Figure 2. The average training time over all models is also shown. Training times are normalized between 0 and 1, where 0 corresponds to the lowest runtime of any model, and 1 corresponds to the largest runtime.
Figure 2 shows that the pre-trained ResNet50 is most capable of generalization based on accuracy and precision. Transfer learning is a well-known method for enhancing model generalization, and should be used when applicable [negative_transfer]. The pre-trained ResNet50 achieves the highest performance with $50$ samples per class (i.e. 500 total augmented samples), scoring an average accuracy of 93% and precision of 95%. Run-time increases linearly with data set size, and the best size in terms of performance and run-time is $n=50$. Even with 10 samples, the pre-trained ResNet is capable of learning enough to achieve 75% accuracy. The other models had a much higher variance across runs, and greatly lagged behind in performance at the levels of data tested.
Interestingly, our results roughly correspond with those originally by [source_dataset] [source_dataset]. In their work, the best performance was found when using $40$ samples per class.
@@ 4.2. Automating Augmentation Selection
For each cGAN model, during training iterations, the discriminator is trained on a batch of size 128, consisting of 64 real samples and 64 synthetic samples generated by the generator. The generator is trained by passing generated images through the frozen discriminator, and using the discriminator output as an error signal. A train--test split of 0.86--0.14 was used to train the discriminator. MPcGAN is designed to generate images of size 128x128x3. The best found settings for the generator and discriminator learning rates were 0.002, and 0.0002, respectively. The generator and discriminator both use Adam optimizer with binary cross entropy as the loss function [Adam]. Dropout is used in the discriminator for regularization, and Leaky ReLus are used for activation functions in the cGAN [Dropout, Leaky].
@@
We have computed Fréchet Inception Distance (FID) and Inception Scores (IS) scores for SYMP data in Table 2. FID and IS are two metrics used to evaluate the quality of synthetic data based on the available ground truth data distribution [FID_IS]. Low FID and high IS correspond to better sample quality. Low FID means the synthetic data better resembles the ground truth. High IS ensures that samples have high variance and clarity. The table also shows the performance before (Regular) and after (Filtered) applying the SYMP-Filter.
@@ 5. Discussion and Conclusions
Microplastics are a problem of growing concern, and current research on the problem is obstructed by limited available data. Our work on GANsemble addresses the issues of small data and class imbalance by automating data augmentation selection and improving synthetic sample generation. We show the data chooser module can be used to automate augmentation selection on a small microplastics data set, and believe it should work in other small data settings as well. MPcGAN is a fairly simple generator and discriminator architecture, and enhancements to the architecture would improve the SYMP data generated. Future work should include implementations of GANsemble with augmentation strategies that do not apply spatial transformations to the image. New augmentations for GANsemble could also include variations of blurring, random erasing [random_erase], or the CutMix algorithm [CutMix].

### v2

@@ Abstract.
*Corresponding author: daniel.platnick@torontomu.ca
@@ 1. Introduction
Oversampling the minority class is a natural remedy for class imbalance [Oversample]. Synthetic Minority Oversampling Technique (SMOTE) is a well known method which produces synthetic samples to fix class imbalances using K-nearest neighbors in the feature space [SMOTE]. heart_spectral_aug heart_spectral_aug performed analysis on different augmentation strategies for convolutional neural network (CNN) classification of medical image spectra, and found that simple strategies such as masking and horizontal flipping can greatly improve model performance. Today, oversampling techniques typically include intelligent data augmentations augmentation_survey. CutMix CutMix introduced a state-of-the-art image augmentation algorithm called CutMix, which maintains the benefits of regional dropout regularization without loss of information. no_best_resampling_approach no_best_resampling_approach showed that cases may vary depending on the nature of the data. Therefore, given imperfect information, the system designer must decide the best augmentation strategy to use. This is the challenge of augmentation selection.
@@
Inspired by recent developments in generative AI GAN; cGAN, we present GANsemble to overcome this challenge and further microplastics research. The proposed method connects ideas from data augmentation and synthetic data generation cGAN_synthdata to provide automatic augmentation strategy selection and enhance model learning on small and imbalanced data sets. The GANsemble data chooser module enables the AI system to manipulate the inputted data through augmentation, generating new augmentation strategies to find the strategy which maximizes model generalization. Specifically, GANsemble performs an n-step factorial search search on inputted augmentation strategies, and employs a deep neural network (DNN) to intelligently select the best strategy $Aug^{*}$. This strategy is then used to oversample the original data set and train a cGAN to generate new synthetic samples. GANsemble can significantly improve the augmentation selection process, by decreasing the need for human intervention. Furthermore, we provide MPcGAN, the first cGAN for generating synthetic microplastics (SYMP) data, and establish baseline FID and IS scores for SYMP data.
@@
Generating synthetic sample data to enhance the training set is an idea increasing in popularity cGAN; diffusion_sample; diffusion_sample; CVAE. This approach has not been applied to microplastics research. To make progress in SYMP data generation, this work applies cGANs to create SYMP data, and baseline SYMP data quality measures are established. MPcGAN is trained to produce class-conditioned SYMP samples that enhance training set size, balance, and variation. Note, this paper builds on the important work of source_dataset source_dataset, which identified the most effective amount of oversampling with augmentation for a small imbalanced microplastics data set. However, Tian et al. introduce data leakage into their analysis by creating augmented data from the training set, and evaluating on the same training set. This is a problem, because in practice, the model will be evaluated on data it has never seen before, not data created from previous training data. Performing a similar study, this research builds on their analysis by using a separate evaluation set to prevent data leakage between train and test sets. Multiple runs are used to increase experimental validity.
@@ 2. Data Set Description
The data set we use was introduced by source_dataset source_dataset, and is the only easily accessible microplastics data set we could find. The raw data and supplementary information on data preparation can be found from these sources source_dataset; data_source2; data_source3. In this section, we review its characteristics and how it was created.
Figure 1. Example of the polar conversion that allows DNNs to treat polymer spectra as images source_dataset.
Laser Directed Infrared (LDIR) and Fourier Transform Infrared (FTIR) spectroscopy convert chemical compounds to spectra data, which has previously been shown amenable for learning by deep neural networks (DNNs) CNN_soil; CNN_spectra; LDIR. The data set used in this study contains 210 samples of polymer spectra across 10 classes: 8 classes of plastic spectra, and 2 non-plastic classes source_dataset. The least represented polymer class, polyacetal, has 5 samples, while the most highly represented class silica has 38. The 210 particle-based spectra were collected from various Dutch water sources such as drinking water, surface water, and sewage, as well as the Agilent Clarity version 1.4.10 software database data_source2. Samples were converted to spectra using the LDIR quantum cascade laser (QCL)-based Agilent chemical imaging system. The LDIR system was used with wavenumbers between 975 and 1800cm-1, and a rotation of $0.5$cm-1 data_source3; source_dataset. An example spectra can be found in the left image of Figure 1.
Tian et al. source_dataset also introduced a visual polarized coordinate transformation which allows microplastic particle spectra to be treated as as images. This enables the use of computer vision models, which have been found to be very useful in practice. The spectra in the microplastics data set are mapped through the polar conversion process as follows: The wavenumbers 975-1800 are mapped from 0°$\degree$ to 360°$\degree$ and the absorption rate becomes the polar coordinate magnitude. The images are then colourized such that colour intensity grows with the absorption value. Figure 1 shows an example of this transformation visually.
@@ 3.1. The Data Chooser Module
The second input is a list of base augmentation strategies to be used. While any set of base data augmentation strategies can be used, for the microplastic data, we use four Gaussian approaches: (1) horizontal flipping and horizontal shifting, (2) blur and rotation, (3) zoom and rotation, and (4) circular and rectangular masking. Base augmentation strategy 1 performs random horizontal flipping and shifting. Random horizontal flips in spectra data have been shown to greatly improve ML classification performance heart_spectral_aug, while shifting ensures there are not too many duplicate samples. Augmentation strategy 2 uses varying levels of blur and random rotations to distort the spectra images. The blur effect is achieved through a Gaussian filter with $\alpha\in[10,23]$ and $\sigma\in[2.8,3.82]$, where $\alpha$ is blur intensity and $\sigma$ is smoothness. Different $\alpha$ and $\sigma$ values were tried and these ranges gave the best results. Augmentation strategy 3 applies random rotations and zoom between $1.0$x and $1.34$x. Augmentation strategy 4 is a Monte Carlo style masking strategy, where circles and rectangles cover different areas of the polymer spectra. Masking has also been effective for augmenting data in heart spectra classification heart_spectral_aug, as it forces the learning algorithm to focus on different regions of the image when masked regions are not present.
@@
In principle, it would be possible to train all $\mathcal{O}(2^{A})$ possible cGANs and thereby directly evaluate the effectiveness of the different combinations. However, even for a small data set, doing so is prohibitively expensive given the time and lack of stability of GAN training. As such, we instead focus on the simpler task of classification using a pre-trained model. That is, we consider each of the $\mathcal{O}(2^{n})$ possible combinations of augmentation methods, and train a classifier $r$ times each, where $r$ is another hyperparameter controlling how extensively the augmentation strategies are explored. The classifier with the best average performance over the $r$ runs is then chosen and used for training the cGAN. For the microplastics data, we use a pre-trained ResNet50 as the classifier for identifying the best augmentation strategy. We refer to the resulting augmentation strategy as $Aug^{*}$.
@@ 4. Experiments and Results
Recall that our data set contains only 210 images. This was split into an unbiased testing set containing 20 samples, with 2 images per class, and a training set containing the remaining 190 samples. The 20 samples were not used in creating the augmented data sets from Table 1 or Figure 2 to prevent data leakage. The 20 samples were later added to help train the cGAN. All code relating to this work can be found here: https://github.com/DanielPlatnick/GANsemble.
@@ 4.1. Identifying An Effective Size For Augmented Data Sets
In our first experiment, we attempt to reproduce the experiments by Tian et al. source_dataset, in which they evaluated the tradeoff between accuracy and training time related to augmented data set size. In the original work, a custom augmentation strategy was used. Our goal is to evaluate if similar conclusions can be made with simpler augmentation strategies and without the data leakage that occurred in the original work.
For simplicity, we tested with horizontal flipping and horizontal shifting. Using this augmentation strategy, we create 10 purely augmented data sets of sizes $100,200,\ldots,1000$, each with balanced classes. We then used this data to train three different DNN models: a CNN with $13,804,510$ trainable parameters, a ResNet50 with $23,528,522$ trainable parameters, and a version of that ResNet50 pre-trained on ImageNet Resnet; ImageNet. Each DNN was trained 5 times and the average results on our test set of 20 images are shown in Figure 2. The average training time over all models is also shown. Training times are normalized between 0 and 1, where 0 corresponds to the lowest runtime of any model, and 1 corresponds to the largest runtime.
Figure 2 shows that the pre-trained ResNet50 is most capable of generalization based on accuracy and precision. Transfer learning is a well-known method for enhancing model generalization, and should be used when applicable negative_transfer. The pre-trained ResNet50 achieves the highest performance with $50$ samples per class (i.e. 500 total augmented samples), scoring an average accuracy of 93% and precision of 95%. Run-time increases linearly with data set size, and the best size in terms of performance and run-time is $n=50$. Even with 10 samples, the pre-trained ResNet is capable of learning enough to achieve 75% accuracy. The other models had a much higher variance across runs, and greatly lagged behind in performance at the levels of data tested.
Interestingly, our results roughly correspond with those originally by source_dataset source_dataset. In their work, the best performance was found when using $40$ samples per class.
@@ 4.2. Automating Augmentation Selection
For each cGAN model, during training iterations, the discriminator is trained on a batch of size 128, consisting of 64 real samples and 64 synthetic samples generated by the generator. The generator is trained by passing generated images through the frozen discriminator, and using the discriminator output as an error signal. A train--test split of 0.86--0.14 was used to train the discriminator. MPcGAN is designed to generate images of size 128x128x3. The best found settings for the generator and discriminator learning rates were 0.002, and 0.0002, respectively. The generator and discriminator both use Adam optimizer with binary cross entropy as the loss function Adam. Dropout is used in the discriminator for regularization, and Leaky ReLus are used for activation functions in the cGAN Dropout; Leaky.
@@
We have computed Fréchet Inception Distance (FID) and Inception Scores (IS) scores for SYMP data in Table 2. FID and IS are two metrics used to evaluate the quality of synthetic data based on the available ground truth data distribution FID_IS. Low FID and high IS correspond to better sample quality. Low FID means the synthetic data better resembles the ground truth. High IS ensures that samples have high variance and clarity. The table also shows the performance before (Regular) and after (Filtered) applying the SYMP-Filter.
@@ 5. Discussion and Conclusions
Microplastics are a problem of growing concern, and current research on the problem is obstructed by limited available data. Our work on GANsemble addresses the issues of small data and class imbalance by automating data augmentation selection and improving synthetic sample generation. We show the data chooser module can be used to automate augmentation selection on a small microplastics data set, and believe it should work in other small data settings as well. MPcGAN is a fairly simple generator and discriminator architecture, and enhancements to the architecture would improve the SYMP data generated. Future work should include implementations of GANsemble with augmentation strategies that do not apply spatial transformations to the image. New augmentations for GANsemble could also include variations of blurring, random erasing random_erase, or the CutMix algorithm CutMix.